If a bag of miniature candy contains about 14 pieces per container, the serving size is 5 pieces, and there are 230 calories per serving. How many pieces of candy are there in the bag? How many calories are in one piece? How many calories are there in the whole bag?
Answer #1

Number of pieces of candy in the bang = 14

Serving size = 5 pieces

Number of calories per serving = 230 calories

Number of calories in one piece = (Number of calories per serving)/(Serving size) = 230/5 = 46 calories

Number of calories in whole bag = (Number of calories in one piece) * (Total Number of pieces) = 46 calories/piece * 14 pieces = 644 calories

Note - Post any doubts/queries in comments section.